# 2D or not 2D? Exploring 3D relativistic magnetic reconnection dynamics with highly accurate numerical simulations

### v1

@@ 1 Introduction
Astrophysical high-energy sources consistently exhibit non-thermal photon spectra, implying the presence of mechanisms that accelerate particles to power-law energy distributions (e.g., Giannios_2005; Celotti_Ghisellini_2008; Zhang_etal_2020). Such sources include flares from pulsars’ and black holes’ magnetospheres (e.g., Lyubarskii_1996; Uzdensky_Spitkovsky_2014; Cerutti_2014; Olmi_PWN_2016; Philippov_Spitkovsky_2018; Nathanail_etal_2022; Ripperda_etal_2022), fast radio or gamma-ray bursts (e.g., Zhang_Yan_2011; McKinney_Uzdensky_2012; Beniamini_Giannios_2017; Mahlmann_etal_2022; Most_Philippov_2023), or in the coronae of black holes’ accretion disks at the base of the relativistic jets of Active Galactic Nuclei and microquasars (e.g., Sironi_Petropoulou_Giannios_2015; Petropoulou_etal_2016). To explain the electromagnetic emission we observe, the aforementioned mechanisms must operate efficiently in the relativistic regime, where the average magnetic energy per particle exceeds the rest mass energy (e.g., Lyutikov_Uzdensky_2003). At macroscopic scales (i.e., scales orders of magnitude larger than the skin depth of the plasma), the dynamics of astrophysical flows is well described within the ideal Magnetohydrodynamics (MHD) approximation, under which the Alfvén theorem (i.e., flux freezing) applies. However, as the flow evolves, localized regions with strong gradients and intense currents may develop. In these regions, non-ideal effects become significant and can no longer be neglected in the energy-momentum balance (e.g., Qian_rHARM_2017; Vourellis_2019; Ripperda_2019a; Ripperda_Bacchini_Philippov_2020; Mattia2023; Nathanail_2025).
@@
In this context, magnetic reconnection has gained popularity over the past few decades as an efficient and universal mechanism that can power the transient, burst-like emissions observed (e.g., Sironi_2025_RecReview, and references therein). It occurs when magnetic field lines of inverse polarity annihilate, rapidly releasing the energy stored in the magnetic field as heat, fast plasma flows, and energetic particles (e.g., Zenitani_2001; Sironi_Spitkovsky_2014; Werner_2016; Sironi_2022; Zhang_2023). For this mechanism to operate, the plasma must depart from ideal behaviour, so that the Alfvén theorem no longer applies locally and magnetic field lines are free to change their topology. In the resistive MHD approximation, the process is regulated by the induction equation where two timescales can be identified: the ideal (alfvénic) one $\tau_{A}=L/c_{A}$, associated with dynamical variations of the magnetic field, and the diffusive one $\tau_{D}=L^{2}/\eta$, associated with the diffusion of the magnetic field. Here $L$ is a characteristic scale of variation of the equilibrium magnetic field, $c_{A}$ is the Alfvén velocity, and $\eta$ is the local magnetic resistivity.
@@
Classical steady-state reconnection models - such as the Sweet-Parker (Sweet_1958; Parker_1957) - fail to predict the highly energetic transient events observed in astrophysical plasmas. These models have therefore been supplanted by descriptions of bursty, tearing-unstable reconnection regions, characterized by the continuous formation and expulsion of multiple secondary islands (i.e., plasmoids Zelenyi_Krasnoselskikh_1979; Bhattacharjee_2009; Uzdensky_etal_2010; Loureiro_Uzdensky_2016). The tendency of current sheets to fragment into chains of magnetic islands or plasmoids was originally formalized by Furth_etal_1963. The predicted reconnection timescale, $\tau_{\mathrm{rec}}\propto\sqrt{\tau_{D}\tau_{A}}$, implies a positive correlation with the Lundquist number, $S=\tau_{D}/\tau_{A}=Lc_{A}/\eta$, which, however, is insufficient to account for the explosive flaring activity observed, given the extremely large values of $S$ typical of relativistic astrophysical plasmas.
@@
Pucci_Velli_2014 presented a theoretical model leading to fast reconnection regimes in the single-fluid resistive MHD approximation. Their work showed that, providing large enough values of the Lundquist number ($S\geq 10^{6}$), Sweet-Parker current sheets are unlikely to form in astrophysical plasmas. Instead, when thin current sheets of inverse aspect ratio $a/L$, where $a$ is the current sheet half width, and $L$ is the current sheet half length, scale as $S^{-1/3}$, a critical transition occurs: the growth rate of the tearing mode instability becomes asymptotically independent of $S$, with modes growing on ideal time scales $\propto\tau_{A}$. The “ideal” tearing instability was verified numerically for the first time by Landi_etal2015, who analyzed the stability of a current sheet undergoing ideal tearing through two-dimensional (2D) compressible MHD simulations, and by Tenerani_2015, who confirmed that in collapsing current sheets, fast reconnection is triggered before a Sweet-Parker-type configuration can form. The fast reconnection regime was also extended to account for visco-resistive (Tenerani_visco_2015), Hall effects (Pucci_Hall_2017; Papini_etal2019), or in double-current sheet systems (Baty_2017_ITMI), always with 2D numerical probes. DelZanna_2016 extended this framework to resistive relativistic MHD, using 2D numerical simulations. To date, studies of reconnection dynamics driven by the ideal tearing mode instability within fully three-dimensional (3D) resistive spacial-relativistic MHD(ResRMHD) frameworks remain unexplored.
@@
Investigating 3D relativistic reconnection under different configurations is crucial to overcome the intrinsic limitations of two-dimensional models and to assess the extent to which such models capture the dynamics occurring in nature. In this paper, we present simulations of 3D relativistic magnetic reconnection triggered by the ideal tearing mode instability by solving the equations of resistive relativistic MHD (Komissarov_2007; DelZanna_2016; Mignone_etal_2019; Ripperda_2019a). While 2D reconnection can be bursty due to the stochastic formation, motion, and merging of plasmoids, these events generally follow a predictable evolution toward a unique final state (e.g., Shibata_Tanuma_2001; Loureiro_2007; Huang_Bhattacharjee_2013; Guo_2016; Hakobyan_2021). In contrast, 3D reconnection exhibits greater variability, with stochastic processes generating a thicker turbulent layer in diverse ways (Cerutti_2014). Following Werner_Uzdensky_2021, we emphasize that reconnection should be understood as the conversion of unreconnected to reconnected magnetic flux, conserving the total flux. In 2D, this conservation is straightforward, but in 3D it is challenging to define and measure it precisely. Rather than adopting a strict definition of reconnection, we focus on assessing the properties of the current layer, such as magnetic and plasma energy, and the layer dynamics. Reconnection in 3D therefore refers broadly to any magnetic-energy-depleting evolution—including flux annihilation—of thin current sheets that could, in principle, undergo 2D-like reconnection, while references to reconnecting flux specifically denote the 2D flux-conserving process.
@@
Resistive relativistic MHD provides a cost-effective framework to capture the impact of magnetic dissipation—occurring on microscopic scales—within the macroscopic evolution of relativistic astrophysical plasmas (e.g. Zanotti_Dumbser_2011; Ripperda_2019b; Bransgrove_Ripperda_2021; Mattia_etal_2024; Grehan_2025). At the same time, the role of physical and numerical resistivity in ResRMHD simulations is also a key question to disentangle (Mattia2023; Grehan_2025; Bugli_etal_2024; Das_2025). A sufficiently high grid resolution is required to ensure that dissipation is set by the imposed physical resistivity rather than by numerical diffusion. This is particularly important for capturing ideal tearing, which develops in low-resistivity regimes and demands sufficient resolution to discretize the smaller dissipation scales at reconnection onset (Puzzoni_2021). In this context, recent advances in $4^{\rm th}$-order accurate finite-volume schemes (Verma_etal2019; Berta_etal_2024; Mignone_Berta_2024) offer significant improvements over traditional $2^{\rm nd}$-order methods. Their lower numerical dissipation enables exploration of reconnection processes with reduced grid resolution requirements while retaining stability and computational efficiency (see, e.g., §4.5 of Mignone_Berta_2024). These capabilities, combined with the power of Graphics Processing Units (GPUs) to reduce computational demands help making fully 3D ResRMHD simulations more accessible. In §2 we describe the working framework and the simulation setups. In §3 we describe the results of the simulations, discussing the results obtained in §4.
@@ 2.1 Equations of ResRMHD
The system of ResRMHD equations consists of a set of conservation laws in the form of hyperbolic differential equations with non-zero source terms (Komissarov_2007; Mignone_etal_2019)
@@
respectively the mass density, momentum density, and total energy density, measured in the laboratory frame. The primitive fluid variables $(\rho,{\mathbf{v}},p)$ are the rest-mass density, the velocity, and the thermal pressure of the fluid, $h$ is the relativistic specific enthalpy function, while $\gamma$ is the Lorentz factor of the flow. In our models, we assume a Taub equation of state (Mignone_2007)
@@
obtained by assuming a simple Ohm’s law with a scalar resistivity $\eta$ in the comoving frame. The term proportional to $\eta^{-1}$ is usually large, due to the extreme conductivity of astrophysical plasmas, it is numerically stiff and requires an implicit treatment (Palenzuela_etal2009; Mignone_etal_2019; Tomei_etal_2020; Mattia2023; Mignone_Berta_2024). In this work we adopt the numerical scheme proposed by Mignone_Berta_2024 to discretize the governing equations.
@@ 2.2 Initial conditions
The parameter $\zeta$, originally introduced by Landi_etal2015, allows one to impose different kinds of initial equilibria: specifically, values of $0<\zeta<1$ interpolate between a purely pressure-balanced configuration for $\zeta=0$, and a force-free equilibrium for $\zeta=1$. The choice $p\propto\rho$ provides an initial uniform temperature across the whole domain $\Theta=\Theta_{0}=p_{0}/\rho_{0}=\sigma_{0}\beta_{0}/2$, as well as the specific upstream enthalpy $h_{0}$, to be retrieved from Eq. (5). The upstream value of the relativistic Alfvén velocity may be written as a function of $\sigma_{0}$ and $\beta_{0}$ as
@@ 2.3 Numerical setup
Our setups are controlled by a set of six parameters, i.e., $\{\sigma_{0},\beta_{0},S,a/L,\varepsilon,\zeta\}$. The parameters $\sigma_{0}$, $\beta_{0}$, $S$, $a/L$, with $L=1$, and $\varepsilon$ are set, respectively, to $\sigma_{0}=10$, $\beta_{0}=0.1$, $S=10^{6}$, $a/L=S^{-1/3}=10^{-2}$, such that the ideal tearing condition is fulfilled, and $\epsilon=10^{-3}$, while $\zeta$ can be either 0 or 1. This choice yields an initial upstream Alfvén velocity of $c_{A}\simeq 0.89$ (in units of $c$), and a resistivity of $\eta\simeq 0.89\times 10^{-6}$. The computational domain spans $x\in[-80a,80a]$, $y,z\in[0,2L]$, with reflective boundary conditions applied along $x$ and periodic boundaries along $y,z$. Based on results provided in Mignone_Berta_2024, simulations were run with a grid resolution of $N_{\rm grid}=\{1024\times 768\times 768\}$ for the 3D cases, and $N_{\rm grid}=\{1024\times 768\}$ for the 2D. The latter simulations have been conducted as reference cases by simply avoiding the dependence on $z$. The final simulation time for all runs is $t=60\tau_{c}$, with $\tau_{c}=c/L$ being the light-crossing time of the sheet’s half length. Table 1 summarizes the models taken under consideration in this article.
@@ 3 Results
In this section, we present the dynamical evolution obtained from our 3D simulations. To assess the impact of dimensionality on the reconnection process, we also compare the properties of the 3D runs with those of corresponding 2D reference simulations. Our models are produced using the numerical scheme described by Mignone_Berta_2024, employing the $5^{\rm th}$-order weighted essentially non-oscillatory (WENOZ) reconstruction of Borges_WENOZ2008. The scheme uses the IMEX SSP(4,3,3) Runge–Kutta time integration method of Pareschi_Russo2005. The implicit portion is solved with an optimized Newton–Broyden multidimensional root finder, while the explicit portion computes edge-centered electromagnetic fields using the multidimensional five-wave Riemann solver of Mignone_etal_2019, and updates zone-centered hydrodynamic variables using a Lax–Friedrichs Riemann solver. All simulations were carried out with the fully parallel, GPU-accelerated version of the PLUTO code (Mignone_PLUTO2007; Rossazza_GPU_2025).
@@ 3.1 Force-free configuration
[…]
@@
At early times, the linear phase of the tearing instability shown by $\left<{B}\right>_{x,\mathrm{RMS}}$ in the upper-left panel of Fig. 1 displays the characteristic exponential growth in both the ff2d and ff3d simulations, ending around $t=8\tau_{c}$ for the 2D case (dotted vertical line). On the other hand, the 3D model evolves more slowly, owing to a redistribution of the energy imparted by the initial perturbation over three dimensions rather than two. The resulting linear phase extends to $t=12\tau_{c}$ (dash-dotted vertical line), attaining a lower value of $\left<{B}\right>_{x,\mathrm{RMS}}$. During the linear evolution, two distinct slopes are observed: the first is associated with the ideal tearing instability, whereas the second corresponds to the linear growth phase of plasmoid coalescence. This is also evident from the top row of Fig. 2, where, at $t=5\tau_{c}$, the growth of the tearing instability in 3D is delayed compared to the 2D case. Towards the end of the linear phase the current sheet undergoes fragmentation and plasmoid formation. This transition is visible both in the emergence of sharp peaks in the integrated current $\left<{|J|}\right$> (upper-right panel of Fig. 1), and in the structures appearing in the middle and bottom rows of Fig. 2. By $t=10\tau_{c}$, the 2D simulation has already entered the non-linear stage, exhibiting a highly structured pattern of current filaments and ripple-like features caused by the non-linear interaction of waves. In contrast, at the same time, the 3D model is still in the linear phase, with the first plasmoids forming within the central current layer. By $t=15\tau_{c}$, the 2D simulation has developed a large, symmetric plasmoid, whereas the 3D simulation produces a more irregular current structure with reduced symmetry and wave interactions, reflecting the additional degrees of freedom available in 3D.
@@
The evolution of the energy diagnostics follow closely. The total electromagnetic energy $\left<{\mathcal{E}}\right>_{\mathrm{EM}}$ (bottom-left panel of Fig. 1) remains nearly constant until the end of the corresponding linear phases, after which $\left<{\mathcal{E}}\right>_{\mathrm{EM}}$ begins to decrease as it is converted—primarily into heat—while the temperature $\left<{\Theta}\right$> (bottom-right panel of Fig. 1) rises. Indeed, the rate of magnetic energy conversion during the early phase ($t\leq 30\tau_{c}$) is lower in 3D than in the 2D case. Differences in the efficiency of magnetic energy release between the two models become even more significant after $t\simeq 30\tau_{c}$. After this stage, both the electromagnetic energy conversion and temperature rising saturate asymptotically in the 2D simulation, with the system reaching a non-reconnecting relaxed state. This corresponds to the end of plasmoid coalescence in the simulation, marked by the formation of a final magnetic island and the establishment of a new quasi-equilibrium (see also the bottom-left panel of Fig. 2). Larger magnetic energy conversion is observed in model ff3d as in 3D, the energy injected into plasmoids can be further dissipated, whereas in 2D the plasmoids remain comparatively stable (Sironi_Spitkovsky_2014; Sironi_Petropoulou_Giannios_2015). The conversion rate remains approximately constant up to $\tau_{c}\simeq 40$, after which it decreases with a corresponding counterpart in the temperature (see also Liu_2011). In model ff3d the dissipation of magnetic energy continues essentially for the whole duration of the simulation.
@@
Both the coalescence instability and the kinking of the flux rope are parasitic instabilities of tearing and promote the emergence of structures in the $z$-direction, favouring the onset of fully 3D turbulence in the computational box. Their subsequent non-linear evolution reduce the coherence length of magnetic and plasma structures along the $z$-direction, opening new channels for field dissipation. In fact, by $t=35\tau_{c}$ (Fig. 3), the flux rope is fully disrupted, with emerging modes and structures propagating in the $z$-direction. Comparing Figs. 33, the centroid of the mode distribution has gradually shifted in the $z$-direction, indicating the slower but persistent growth of perturbations along this axis. This disruption promotes the onset of self-generated turbulence and the thickening of the reconnection layer. This result is consistent with first-principle kinetic models which likewise exhibit a chaotic, 3D turbulent evolution (see, e.g., Werner_Uzdensky_2021, and references therein).

### v2

@@ 1 Introduction
Astrophysical high-energy sources consistently exhibit non-thermal photon spectra, implying the presence of mechanisms that accelerate particles to power-law energy distributions (e.g., Giannios, 2005; Celotti and Ghisellini, 2008; Zhang et al., 2020). Such sources include flares from pulsars’ and black holes’ magnetospheres (e.g., Lyubarskii, 1996; Baty et al., 2013; Uzdensky and Spitkovsky, 2014; Cerutti et al., 2014; Olmi et al., 2016; Philippov and Spitkovsky, 2018; Nathanail et al., 2022; Ripperda et al., 2022b), fast radio or gamma-ray bursts (e.g., Zhang and Yan, 2011; McKinney and Uzdensky, 2012; Beniamini and Giannios, 2017; Mahlmann et al., 2022; Most and Philippov, 2023), or in the coronae of black holes’ accretion disks at the base of the relativistic jets of Active Galactic Nuclei and microquasars (e.g., Sironi et al., 2015; Petropoulou et al., 2016). To explain the electromagnetic emission we observe, the aforementioned mechanisms must operate efficiently in the relativistic regime, where the average magnetic energy per particle exceeds the rest mass energy (e.g., Lyutikov and Uzdensky, 2003). At macroscopic scales (i.e., scales orders of magnitude larger than the skin depth of the plasma), the dynamics of astrophysical flows is well described within the ideal Magnetohydrodynamics (MHD) approximation, under which the Alfvén theorem (i.e., flux freezing) applies. However, as the flow evolves, localized regions with strong gradients and intense currents may develop. In these regions, non-ideal effects become significant and can no longer be neglected in the energy-momentum balance (e.g., Qian et al., 2017; Vourellis et al., 2019; Ripperda et al., 2019a, 2020; Mattia et al., 2023; Nathanail et al., 2025).
@@
In this context, magnetic reconnection has gained popularity over the past few decades as an efficient and universal mechanism that can power the transient, burst-like emissions observed (e.g., Sironi et al., 2025, and references therein). It occurs when magnetic field lines of inverse polarity annihilate, rapidly releasing the energy stored in the magnetic field as heat, fast plasma flows, and energetic particles (e.g., Zenitani and Hoshino, 2001; Sironi and Spitkovsky, 2014; Werner et al., 2016; Barkov and Komissarov, 2016; Sironi, 2022; Zhang et al., 2023). For this mechanism to operate, the plasma must depart from ideal behaviour, so that the Alfvén theorem no longer applies locally and magnetic field lines are free to change their topology. In the resistive MHD approximation, the process is regulated by the induction equation where two timescales can be identified: the ideal (alfvénic) one $\tau_{A}=L/c_{A}$, associated with dynamical variations of the magnetic field, and the diffusive one $\tau_{D}=L^{2}/\eta$, associated with the diffusion of the magnetic field. Here $L$ is a characteristic scale of variation of the equilibrium magnetic field, $c_{A}$ is the Alfvén velocity, and $\eta$ is the local magnetic resistivity.
@@
Classical steady-state reconnection models - such as the Sweet-Parker (Sweet, 1958; Parker, 1957) - fail to predict the highly energetic transient events observed in astrophysical plasmas. These models have therefore been supplanted by descriptions of bursty, tearing-unstable reconnection regions, characterized by the continuous formation and expulsion of multiple secondary islands (i.e., plasmoids Zelenyi and Krasnoselskikh, 1979; Bhattacharjee et al., 2009; Uzdensky et al., 2010; Loureiro and Uzdensky, 2016). The tendency of current sheets to fragment into chains of magnetic islands or plasmoids was originally formalized by Furth et al. (1963). The predicted reconnection timescale, $\tau_{\mathrm{rec}}\propto\sqrt{\tau_{D}\tau_{A}}$, implies a positive correlation with the Lundquist number, $S=\tau_{D}/\tau_{A}=Lc_{A}/\eta$, which, however, is insufficient to account for the explosive flaring activity observed, given the extremely large values of $S$ typical of relativistic astrophysical plasmas.
@@
Pucci and Velli (2014) presented a theoretical model leading to fast reconnection regimes in the single-fluid resistive MHD approximation. Their work showed that, providing large enough values of the Lundquist number ($S\geq 10^{6}$), Sweet-Parker current sheets are unlikely to form in astrophysical plasmas. Instead, when thin current sheets of inverse aspect ratio $a/L$, where $a$ is the current sheet half width, and $L$ is the current sheet half length, scale as $S^{-1/3}$, a critical transition occurs: the growth rate of the tearing mode instability becomes asymptotically independent of $S$, with modes growing on ideal time scales $\propto\tau_{A}$. The “ideal” tearing instability was verified numerically for the first time by Landi et al. (2015), who analyzed the stability of a current sheet undergoing ideal tearing through two-dimensional (2D) compressible MHD simulations, and by Tenerani et al. (2015b), who confirmed that in collapsing current sheets, fast reconnection is triggered before a Sweet-Parker-type configuration can form. The fast reconnection regime was also extended to account for visco-resistive (Tenerani et al., 2015a), Hall effects (Pucci et al., 2017; Papini et al., 2019), or in double-current sheet systems (Komissarov et al., 2007; Baty, 2017), always with 2D numerical probes. Del Zanna et al. (2016) extended this framework to resistive relativistic MHD, using 2D numerical simulations. To date, studies of reconnection dynamics driven by the ideal tearing mode instability within fully three-dimensional (3D) resistive spacial-relativistic MHD(ResRMHD) frameworks remain unexplored.
@@
Investigating 3D relativistic reconnection under different configurations is crucial to overcome the intrinsic limitations of two-dimensional models and to assess the extent to which such models capture the dynamics occurring in nature. In this paper, we present simulations of 3D relativistic magnetic reconnection triggered by the ideal tearing mode instability by solving the equations of resistive relativistic MHD (Komissarov, 2007; Del Zanna et al., 2016; Mignone et al., 2019; Ripperda et al., 2019a). While 2D reconnection can be bursty due to the stochastic formation, motion, and merging of plasmoids, these events generally follow a predictable evolution toward a unique final state (e.g., Shibata and Tanuma, 2001; Loureiro et al., 2007; Komissarov et al., 2007; Huang and Bhattacharjee, 2013; Guo et al., 2016; Hakobyan et al., 2021). In contrast, 3D reconnection exhibits greater variability, with stochastic processes generating a thicker turbulent layer in diverse ways (Cerutti et al., 2014). Following Werner and Uzdensky (2021), we emphasize that reconnection should be understood as the conversion of unreconnected to reconnected magnetic flux, conserving the total flux. In 2D, this conservation is straightforward, but in 3D it is challenging to define and measure it precisely. Rather than adopting a strict definition of reconnection, we focus on assessing the properties of the current layer, such as magnetic and plasma energy, and the layer dynamics. Reconnection in 3D therefore refers broadly to any magnetic-energy-depleting evolution—including flux annihilation—of thin current sheets that could, in principle, undergo 2D-like reconnection, while references to reconnecting flux specifically denote the 2D flux-conserving process.
@@
Resistive relativistic MHD provides a cost-effective framework to capture the impact of magnetic dissipation—occurring on microscopic scales—within the macroscopic evolution of relativistic astrophysical plasmas (e.g. Zanotti and Dumbser, 2011; Ripperda et al., 2019b; Bransgrove et al., 2021; Mattia et al., 2024; Grehan et al., 2025). At the same time, the role of physical and numerical resistivity in ResRMHD simulations is also a key question to disentangle (Mattia et al., 2023; Grehan et al., 2025; Bugli et al., 2025; Das et al., 2025; Ripperda et al., 2026). A sufficiently high grid resolution is required to ensure that dissipation is set by the imposed physical resistivity rather than by numerical diffusion. This is particularly important for capturing ideal tearing, which develops in low-resistivity regimes and demands sufficient resolution to discretize the smaller dissipation scales at reconnection onset (Puzzoni et al., 2021).
In this context, recent advances in $4^{\rm th}$-order accurate finite-volume schemes (Verma et al., 2019; Berta et al., 2024; Mignone et al., 2024) offer significant improvements over traditional $2^{\rm nd}$-order methods. Their lower numerical dissipation enables exploration of reconnection processes with reduced grid resolution requirements while retaining stability and computational efficiency (see, e.g., §4.5 of Mignone et al., 2024). These capabilities, combined with the power of Graphics Processing Units (GPUs) to reduce computational demands help making fully 3D ResRMHD simulations more accessible. In §2 we describe the working framework and the simulation setups. In §3 we describe the results of the simulations, discussing the results obtained in §4.
@@ 2.1 Equations of ResRMHD
The system of ResRMHD equations consists of a set of conservation laws in the form of hyperbolic differential equations with non-zero source terms (Komissarov, 2007; Mignone et al., 2019)
@@
respectively the mass density, momentum density, and total energy density, measured in the laboratory frame. The primitive fluid variables $(\rho,{\mathbf{v}},p)$ are the rest-mass density, the velocity, and the thermal pressure of the fluid, $h$ is the relativistic specific enthalpy function, while $\gamma$ is the Lorentz factor of the flow. In our models, we assume a Taub equation of state (Mignone, 2007)
@@
obtained by assuming a simple Ohm’s law with a scalar resistivity $\eta$ in the comoving frame. The term proportional to $\eta^{-1}$ is usually large, due to the extreme conductivity of astrophysical plasmas, it is numerically stiff and requires an implicit treatment (Palenzuela et al., 2009; Mignone et al., 2019; Tomei et al., 2020; Mattia et al., 2023; Mignone et al., 2024). In this work we adopt the numerical scheme proposed by Mignone et al. (2024) to discretize the governing equations.
@@ 2.2 Initial conditions
The parameter $\zeta$, originally introduced by Landi et al. (2015), allows one to impose different kinds of initial equilibria: specifically, values of $0<\zeta<1$ interpolate between a purely pressure-balanced configuration for $\zeta=0$, and a force-free equilibrium for $\zeta=1$. The choice $p\propto\rho$ provides an initial uniform temperature across the whole domain $\Theta=\Theta_{0}=p_{0}/\rho_{0}=\sigma_{0}\beta_{0}/2$, as well as the specific upstream enthalpy $h_{0}$, to be retrieved from Eq. (5). The upstream value of the relativistic Alfvén velocity may be written as a function of $\sigma_{0}$ and $\beta_{0}$ as
@@ 2.3 Numerical setup
Our setups are controlled by a set of six parameters, i.e., $\{\sigma_{0},\beta_{0},S,a/L,\varepsilon,\zeta\}$. The parameters $\sigma_{0}$, $\beta_{0}$, $S$, $a/L$, with $L=1$, and $\varepsilon$ are set, respectively, to $\sigma_{0}=10$, $\beta_{0}=0.1$, $S=10^{6}$, $a/L=S^{-1/3}=10^{-2}$, such that the ideal tearing condition is fulfilled, and $\epsilon=10^{-3}$, while $\zeta$ can be either 0 or 1. This choice yields an initial upstream Alfvén velocity of $c_{A}\simeq 0.89$ (in units of $c$), and a resistivity of $\eta\simeq 0.89\times 10^{-6}$. The computational domain spans $x\in[-80a,80a]$, $y,z\in[0,2L]$, with reflective boundary conditions applied along $x$ and periodic boundaries along $y,z$. Based on results provided in Mignone et al. (2024), simulations were run with a grid resolution of $N_{\rm grid}=\{1024\times 768\times 768\}$ for the 3D cases, and $N_{\rm grid}=\{1024\times 768\}$ for the 2D. The latter simulations have been conducted as reference cases by simply avoiding the dependence on $z$. The final simulation time for all runs is $t=60\tau_{c}$, with $\tau_{c}=c/L$ being the light-crossing time of the sheet’s half length. Table 1 summarizes the models taken under consideration in this article.
@@ 3 Results
In this section, we present the dynamical evolution obtained from our 3D simulations. To assess the impact of dimensionality on the reconnection process, we also compare the properties of the 3D runs with those of corresponding 2D reference simulations. Our models are produced using the numerical scheme described by Mignone et al. (2024), employing the $5^{\rm th}$-order weighted essentially non-oscillatory (WENOZ) reconstruction of Borges et al. (2008). The scheme uses the IMEX SSP(4,3,3) Runge–Kutta time integration method of Pareschi and Russo (2005). The implicit portion is solved with an optimized Newton–Broyden multidimensional root finder, while the explicit portion computes edge-centered electromagnetic fields using the multidimensional five-wave Riemann solver of Mignone et al. (2019), and updates zone-centered hydrodynamic variables using a Lax–Friedrichs Riemann solver. All simulations were carried out with the fully parallel, GPU-accelerated version of the PLUTO code (Mignone et al., 2007; Rossazza et al., 2025).
@@ 3.1 Force-free configuration
Alongside are plotted several other relevant volume-integrated quantities: the norm of $|\nabla\times\mathbf{B}|$ used as a proxy for the non-relativistic electric current $\left<{|J|}\right$>; the electromagnetic energy normalized by its initial value $\left<{\mathcal{E}}\right>_{\mathrm{EM}}=\left(\left<{E^{2}}\right>+\left<{B^{2}}\right>\right)/\left<{B_{0}^{2}}\right$>; and the average temperature $\left<{\Theta}\right>=\left<{p/\rho}\right$>.
@@
At early times, the linear phase of the tearing instability shown by $\left<{B}\right>_{x,\mathrm{RMS}}$ in the upper-left panel of Fig. 1 displays the characteristic exponential growth in both the ff2d and ff3d simulations, ending around $t=8\tau_{c}$ for the 2D case (dotted vertical line). On the other hand, the 3D model evolves more slowly, owing to a redistribution of the energy imparted by the initial perturbation over three dimensions rather than two. The resulting linear phase extends to $t=12\tau_{c}$ (dash-dotted vertical line), attaining a lower value of $\left<{B}\right>_{x,\mathrm{RMS}}$. During the linear evolution, two distinct slopes are observed: the first is associated with the ideal tearing instability, whereas the second corresponds to the linear growth phase of plasmoid coalescence. This is also evident from the top row of Fig. 2, where, at $t=5\tau_{c}$, the growth of the tearing instability in 3D is delayed compared to the 2D case. Towards the end of the linear phase the current sheet undergoes fragmentation and plasmoid formation. This transition is visible both in the emergence of sharp peaks in the integrated current $\left<{|J|}\right$> (upper-right panel of Fig. 1), and in the structures appearing in the middle and bottom rows of Fig. 2. By $t=10\tau_{c}$, the 2D simulation has entered the non-linear stage, exhibiting a highly structured pattern of current filaments and ripple-like features caused by the non-linear interaction of waves. At the same time, the 3D model is still in the linear phase, with the first plasmoids forming within the central current layer. By $t=15\tau_{c}$, the 2D simulation has developed a large, symmetric plasmoid, whereas the 3D simulation produces a more irregular current structure with reduced symmetry and wave interactions, reflecting the additional degrees of freedom available in 3D.
@@
The evolution of the energy diagnostics follow closely. The total electromagnetic energy $\left<{\mathcal{E}}\right>_{\mathrm{EM}}$ (bottom-left panel of Fig. 1) remains nearly constant until the end of the corresponding linear phases, after which $\left<{\mathcal{E}}\right>_{\mathrm{EM}}$ begins to decrease as it is converted—primarily into heat—while the temperature $\left<{\Theta}\right$> (bottom-right panel of Fig. 1) rises. Indeed, the rate of magnetic energy conversion during the early phase ($t\leq 30\tau_{c}$) is lower in 3D than in the 2D case. Differences in the efficiency of magnetic energy release between the two models become even more significant after $t\simeq 30\tau_{c}$. After this stage, both the electromagnetic energy conversion and temperature rising saturate asymptotically in the 2D simulation, with the system reaching a non-reconnecting relaxed state. This corresponds to the end of plasmoid coalescence in the simulation, marked by the formation of a final magnetic island and the establishment of a new quasi-equilibrium (see also the bottom-left panel of Fig. 2). Larger magnetic energy conversion is observed in model ff3d as in 3D, the energy injected into plasmoids can be further dissipated, whereas in 2D the plasmoids remain comparatively stable (Sironi and Spitkovsky, 2014; Sironi et al., 2015). The conversion rate remains approximately constant up to $\tau_{c}\simeq 40$, after which it decreases with a corresponding counterpart in the temperature (see also Liu et al., 2011). In model ff3d the dissipation of magnetic energy continues essentially for the whole duration of the simulation.
@@
Both the coalescence instability and the kinking of the flux rope are parasitic instabilities of tearing and promote the emergence of structures in the $z$-direction, favouring the onset of fully 3D turbulence in the computational box. Their subsequent non-linear evolution reduce the coherence length of magnetic and plasma structures along the $z$-direction, opening new channels for field dissipation. In fact, by $t=35\tau_{c}$ (Fig. 3), the flux rope is fully disrupted, with emerging modes and structures propagating in the $z$-direction. Comparing Figs. 33, the centroid of the mode distribution has gradually shifted in the $z$-direction, indicating the slower but persistent growth of perturbations along this axis. This disruption promotes the onset of self-generated turbulence and the thickening of the reconnection layer. This result is consistent with first-principle kinetic models which likewise exhibit a chaotic, 3D turbulent evolution (see, e.g., Werner and Uzdensky, 2021, and references therein).